# RoleLLM: Benchmarking, Eliciting, and Enhancing Role-Playing Abilities of Large Language Models

### v2

@@ 5.1 Experiment Setup
The LLaMA-7B model (Touvron et al.,, 2023) is fine-tuned on both RoleBench-general-en and RoleBench-specific-en utilizing LoRA tuning (Hu et al.,, 2022). Similarly, the ChatGLM2-6B111111https://github.com/THUDM/ChatGLM2-6B model (Zeng et al.,, 2022; Du et al.,, 2022) is fine-tuned on RoleBench-general-zh and RoleBench-specific-zh employing the same technique. It is worth noting that LLaMA-7B is only a pre-trained model, and ChatGLM2-6B is post-trained with enhanced instruction-following and dialogue capabilities. Please refer to Appendix F for more details.
@@
RoleGPT is a strong baseline for both Chinese and English roles (GPT-4). For English roles, LLaMA-7B and its instruction-tuned variants, Vicuna-13B (Chiang et al.,, 2023) and Alpaca-7B (Taori et al.,, 2023), are as baselines. For Chinese roles, ChatGLM2-6B is a baseline. We also incorporate models trained on script data in a multi-turn conversation mode as additional baselines.
@@
We employ Rouge-L (Lin,, 2004) to measure the overlap between model predictions and ground truths. The ground truths comprise three categories: (1) Raw ground-truths of general instructions without role-playing (RAW); (2) Customized general instruction responses with role-playing from RoleBench-general (CUS); (3) Role-specific instruction responses from RoleBench-specific (SPE). RAW assesses the model’s response accuracy to instructions. CUS gauges the model’s ability to mimic the speaking style associated with a particular role. SPE tests the model’s role-specific knowledge and memories. Following AlpacaEval (Li et al., 2023c, ), we also use GPT as an evaluator, given the proven reliability of GPT evaluators (Fu et al.,, 2023; Gilardi et al.,, 2023; Zheng et al.,, 2023). GPT evaluator prompts, with minor modifications from AlpacaEval’s121212https://github.com/tatsu-lab/alpaca_eval/blob/main/src/alpaca_eval/evaluators_configs/alpaca_eval_gpt4/alpaca_eval.txt, are detailed in Appendix K.5. These prompts aid in sample comparison and ranking to obtain a win rate or an average ranking. Our train-test set splitting strategy focuses on two dimensions: (1) Instruction-based splits for assessing instruction generalization; and (2) Role-based splits for assessing role generalization, limited to English. Please refer to Appendix G for more details.
@@ 5.2 Main Experiments
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 0.5764 | 0.5322 | 0.3229 | 0.4772 |
| LLaMA | 0.1291 | 0.1232 | 0.2553 | 0.1692 |
| LLaMA-script | 0.0832 | 0.0509 | 0.1081 | 0.0807 |
| Alpaca | 0.2423 | 0.3529 | 0.2695 | 0.2882 |
| Vicuna | 0.2101 | 0.2550 | 0.2913 | 0.2521 |
| RoleLLaMA | 0.3294 | 0.3755 | 0.3814 | 0.3621 |
| Model | Win Rate |
| Vicuna | 29.26 |
| Alpaca | 29.30 |
| RoleLLaMA | 55.82 |
@@
Tables 3(a) and 3(b) present the Rouge-L and GPT scores respectively, for the evaluation of instruction generalization abilities. For the GPT-3.5 evaluation, each model is compared with our RoleGPT to determine a win rate, signifying the frequency at which a model is favored over RoleGPT by GPT-3.5. We observe that RoleLLaMA manifests a considerable enhancement in role-playing performance compared to instruction-tuned baselines and the vanilla base model, in terms of speaking style imitation (CUS), response accuracy (RAW), and role-specific knowledge (SPE). Moreover, LLaMA-script directly tuned on original conversation data in scripts, even underperforms base LLaMA, which shows the imperative need for denoising of script data. The comparison between Alpaca/Vicuna and LLaMA indicates the efficacy of general-purpose instruction tuning in role-playing. However, the discrepancy between RoleLLaMA and Alpaca/Vicuna underscores the necessity for LLMs to enhance their role-playing capabilities beyond mere general-purpose instruction-following. Additionally, despite RoleLLaMA trailing behind RoleGPT in terms of speaking style imitation (CUS) and response accuracy (RAW)131313Note that drawing a direct comparison between RoleLLaMA and RoleGPT may not be equitable, given that RoleGPT is closed-source and the specifics of its model size and instruction tuning data remain undisclosed., it surpasses RoleGPT in role-specific knowledge, indicating the efficacy of our Context-Instruct technique.
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 0.5368 | 0.5748 | 0.2478 | 0.4531 |
| ChatGLM2 | 0.3938 | 0.506 | 0.3099 | 0.4032 |
| ChatGLM2-script | 0.1401 | 0.3066 | 0.0915 | 0.1794 |
| RoleGLM | 0.5049 | 0.5255 | 0.3406 | 0.4570 |
| Model | Win Rate |
| ChatGLM2 | 24.17 |
| RoleGLM | 36.37 |
@@
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 0.6020 | 0.5322 | 0.2985 | 0.4776 |
| LLaMA | 0.1316 | 0.1232 | 0.2551 | 0.2238 |
| Alpaca | 0.2315 | 0.3529 | 0.2587 | 0.3015 |
| Vicuna | 0.2082 | 0.2550 | 0.2775 | 0.2842 |
| RoleLLaMA | 0.4126 | 0.4107 | 0.2568 | 0.3600 |
| Model | Win Rate |
| Alpaca | 16.96 |
| Vicuna | 37.65 |
| RoleLLaMA | 49.51 |
@@
In Table 5(a) and Table 5(b), considering 10 held-out unseen roles, we observe that RoleLLaMA exhibits performance increase in imitating speaking style (CUS) and maintaining accuracy (RAW). However, with respect to role-specific knowledge (SPE), there is not a noticeable improvement compared to the baselines, which is intuitive since models with no prior knowledge of an unseen role would naturally exhibit confusion regarding it.
@@ 5.3 Ablation Study
| RoleGPT (zsp) | 9.85 | 2.31 |
| RoleGPT (fsp) | 30.19 | 1.92 |
| RoleGPT (fsd) | 61.79 | 1.49 |
@@
We employ GPT-3.5 to evaluate three role prompting methods in §3.2, namely, zero-shot prompt engineering (zsp), few-shot prompt engineering (fsp), and few-shot dialogue engineering (fsd). The evaluation is based on RoleBench-specific-zh in Table 6. The Win Rate is the frequency with which a method is ranked first by GPT-3.5 among the three and Avg. Ranking refers to the average ranking position of the method. We observe that few-shot dialogue engineering significantly outperforms few-shot prompt engineering, and both few-shot approaches excel over the zero-shot approach. These findings underscore the significance of dialogue-mode input formatting and retrieval augmentation for GPT.
| Method | RoleLLaMA (reaug) | RoleLLaMA (sys) | RoleGLM (reaug) | RoleGLM (sys) | RoleGPT (sys) | RoleGPT (reaug) |
| SPE | 0.3671 | 0.3814 | 0.2525 | 0.3406 | 23.36 | 23.59 |
@@
We compare two role customization approaches: zero-shot system instruction (sys, §3.4) v.s. few-shot retrieval augmentation (reaug). Retrieval-augmentation-based role customization is akin to the way mentioned in §3.2, we insert retrieved dialogue pairs from profiles into the input, formatted according to the chat markup language specific to RoleLLaMA and RoleGLM. During both fine-tuning and inference, the input combines system instruction and these retrieved in-context demonstrations. Table 7 shows the superior performance of the system-instruction-based approach over the retrieval-augmentation-based approach for RoleLLaMA and RoleGLM, leading to a higher context efficiency. We suppose that the retrieved in-context examples from profiles are noisy and sparse, which leads to the distraction of the relatively small LLMs (e.g., RoleGLM and RoleLLaMA) and degrades their performance. While larger models such as RoleGPT exhibit greater robustness to noisy and sparse information, leading to performance invariability and even performance increase (Li et al.,, 2022; Shi et al.,, 2023).
| Method | SPE |
| RoleLLaMA (w/o c-inst) | 0.2140 |
| RoleLLaMA (reaug, w/o c-inst) | 0.1914 |
| RoleLLaMA (w/ c-inst) | 0.3814 |
| Method | SPE |
| RoleLLaMA (w/o c-inst) | 0.2809 |
| RoleLLaMA (reaug, w/o c-inst) | 0.1866 |
| RoleLLaMA (w/ c-inst) | 0.3504 |
Effect of Context-Instruct.
To inject role-specific knowledge, a typical approach is to employ retrieval-augmentation role customization strategy (c.f., §3.2 & §5.3). In Table 8, we explore the effect of Context-Instruct to inject role-specific knowledge, comparing it with retrieval augmentation and the absence of both techniques. We observe that Context-Instruct has the capability to substantially augment the models’ role-specific knowledge, and retrieval augmentation may lead to distraction and a lack of robustness due to the noisy nature of the retrieval source discussed in §5.3.
@@
In Figure 6, we analyze the scaling law of role-playing of RoleLLaMA with different sizes (i.e., 7B, 13B, 33B), and observe that as the model size increases, the results of all metrics (i.e., RAW, CUS, SPE) are better.
We refer readers to Appendix H for more analysis (e.g., ablation studies on data mixing strategies).
@@ 6 Conclusion
We introduce RoleLLM, a framework to benchmark, elicit, and enhance role-playing in LLMs. We construct 100 role profiles, propose RoleGPT for speaking style imitation, and Context-Instruct for role-specific knowledge extraction. Through these methods, we create RoleBench, the first fine-grained benchmark and open-source instruction tuning data for role-playing. We obtain RoleLLaMA and RoleGLM by role-conditioned instruction tuning on RoleBench, which demonstrate strong role-playing performance and are even comparable with RoleGPT (using GPT-4).
Contributions
The contributions are listed as follows: (1) Role Collection and Profile Construction: Zekun Moore Wang, Zhongyuan Peng, Yuhan Wu, Junran Peng; (2) RoleBench Construction and Analysis: Zekun Moore Wang (RoleBench-general, quality inspection), Haoran Que (RoleBench-specific), Yuhan Wu (analysis, quality inspection), Zhongyuan Peng (quality inspection), Jiaheng Liu (quality inspection), Hongcheng Guo (quality inspection), Junran Peng (quality inspection), Zehao Ni (quality inspection), Jian Yang (quality inspection); (3) Experiments: Zekun Moore Wang (design, RoleGPT, GPT evaluation), Haoran Que (RoleLLaMA), Zhongyuan Peng (RoleGLM); (4) Context-Instruct: Zekun Moore Wang (design), Haoran Que (implementation); (5) System Demonstration: Zhongyuan Peng, Ruitong Gan; (6) Paper Writing: Zekun Moore Wang, Jiaheng Liu, Haoran Que, Jie Fu; (7) Supervision: Jiaheng Liu, Junran Peng, Zekun Moore Wang; (8) Advisors: Jiaheng Liu, Junran Peng, Jie Fu, Wanli Ouyang, Wangchunshu Zhou, Zhaoxiang Zhang, Ke Xu, Man Zhang, Stephen W. Huang. Junran Peng and Jiaheng Liu initialized the project. Zekun Moore Wang, Jiaheng Liu, and Junran Peng co-led the project. Zekun Moore Wang designed the framework and experiments, and wrote the paper. All contributors engaged in the discussion.
Appendix A A Ethics Statement
Since role-playing may lead to jailbreaking (Liu et al.,, 2023), misuse of RoleLLM could generate sensitive or harmful content, reflecting the violent, explicit, or discriminatory traits of certain characters. This work and its assets are strictly for research purposes, opposing any malicious usage. It’s recommended to employ moderation and filtering mechanisms to curb adverse content dissemination. The responsibility lies with researchers and users to ensure ethical application of RoleLLM.
Appendix B B Non-Cherry-Picked Demonstrations

### v3

@@ 4.1 Experiment Setup
The LLaMA-7B model (Touvron et al., 2023) is fine-tuned on both RoleBench-general-en and RoleBench-specific-en utilizing LoRA tuning (Hu et al., 2022). Similarly, the ChatGLM2-6B111111https://github.com/THUDM/ChatGLM2-6B model (Zeng et al., 2022; Du et al., 2022) is fine-tuned on RoleBench-general-zh and RoleBench-specific-zh employing the same technique. It is worth noting that LLaMA-7B is only a pre-trained model, and ChatGLM2-6B is post-trained with enhanced instruction-following and dialogue capabilities. For simplicity, unless specifically noted, RoleLLaMA by default refers to the role-playing models trained from LLaMA-7B. Please refer to Appendix E for more details.
@@
RoleGPT is a strong baseline for both Chinese and English roles (GPT-4). For English roles, LLaMA-7B and its instruction-tuned variants, Vicuna-13B (Chiang et al., 2023) and Alpaca-7B (Taori et al., 2023), are as baselines. For Chinese roles, ChatGLM2-6B and Yi-6B-Chat121212https://huggingface.co/01-ai/Yi-6B-Chat are baselines. Character.AI131313https://beta.character.ai/ and ChatPLUG Tian et al. (2023) are two baselines specifically designed for role-playing. We also use models trained on script data in a multi-turn conversation mode as additional baselines.
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 57.6 | 53.2 | 32.3 | 47.7 |
| LLaMA | 12.9 | 12.3 | 25.5 | 16.9 |
| LLaMA-script | 8.3 | 5.1 | 10.8 | 8.1 |
| Alpaca | 24.2 | 35.3 | 27.0 | 28.8 |
| Vicuna | 21.0 | 25.5 | 29.1 | 25.2 |
| ChatPLUG | 24.0 | 34.7 | 25.8 | 28.2 |
| Character.AI | 41.9 | 45.7 | 30.3 | 39.3 |
| LLaMA-2-7B-Chat | 16.2 | 26.2 | 21.4 | 21.2 |
| RoleLLaMA-7B | 32.9 | 37.6 | 38.1 | 36.2 |
| RoleLLaMA2 (13B-Chat) | 37.5 | 47.9 | 48.8 | 44.7 |
| Model | GPT-4 | Human |
| Vicuna | 32.0 | 23.4 |
| Alpaca | 16.0 | 20.0 |
| ChatPLUG | 3.8 | 16.4 |
| Character.AI | 31.4 | 30.2 |
| RoleLLaMA-7B | 55.8 | 52.0 |
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 53.7 | 57.5 | 24.8 | 45.3 |
| ChatGLM2 | 39.4 | 50.6 | 31.0 | 40.3 |
| ChatGLM2-script | 14.0 | 30.7 | 9.1 | 17.9 |
| ChatPLUG | 38.9 | 61.4 | 31.0 | 43.8 |
| Character.AI | 42.0 | 55.8 | 28.7 | 42.2 |
| Yi-6B-Chat | 40.6 | 56.5 | 26.9 | 41.3 |
| RoleGLM | 50.5 | 52.6 | 34.1 | 45.7 |
| Model | GPT-4 | Human |
| ChatGLM2 | 24.2 | 19.6 |
| ChatPLUG | 28.9 | 19.9 |
| Character.AI | 28.2 | 19.0 |
| RoleGLM | 36.4 | 52.4 |
| Model | CUS | RAW | SPE | avg. |
| RoleGPT | 60.2 | 53.2 | 29.9 | 47.8 |
| LLaMA | 13.2 | 12.3 | 25.5 | 22.4 |
| Alpaca | 23.2 | 35.3 | 25.9 | 30.2 |
| Vicuna | 20.8 | 25.5 | 27.8 | 28.4 |
| ChatPLUG | 28.7 | 34.7 | 25.0 | 29.5 |
| LLaMA-2-7B-Chat | 20.8 | 26.2 | 20.2 | 22.4 |
| RoleLLaMA-7B | 41.3 | 41.1 | 25.7 | 36.0 |
| Model | GPT-4 | Human |
| Alpaca | 12.0 | 28.2 |
| Vicuna | 31.0 | 32.4 |
| ChatPLUG | 8.8 | 12.0 |
| RoleLLaMA-7B | 64.5 | 56.1 |
@@
We employ Rouge-L (Lin, 2004) to measure the overlap between model predictions and ground truths. The ground truths comprise three categories: (1) Raw ground-truths of general instructions without role-playing (RAW); (2) Customized general instruction responses with role-playing from RoleBench-general (CUS); (3) Role-specific instruction responses from RoleBench-specific (SPE). RAW assesses the model’s response accuracy to instructions. CUS gauges the model’s ability to mimic the speaking style associated with a particular role. SPE tests the model’s role-specific knowledge and memories. Following AlpacaEval (Li et al., 2023c), we also use GPT as an evaluator, given the proven reliability of GPT evaluators (Fu et al., 2023; Gilardi et al., 2023; Zheng et al., 2023). GPT evaluator prompts, with minor modifications from AlpacaEval’s151515https://github.com/tatsu-lab/alpaca_eval/blob/main/src/alpaca_eval/evaluators_configs/alpaca_eval_gpt4/alpaca_eval.txt, are detailed in Appendix I.5. These prompts aid in sample comparison and ranking to obtain a win rate or an average ranking. Our train-test set splitting strategy focuses on two dimensions: (1) Instruction-based splits for assessing instruction generalization; and (2) Role-based splits for assessing role generalization, limited to English. Note that we test Character.AI on a subset with 500 instructions in the test split as it doesn’t offer an API access and we have to manually operate on its website and collect results. Furthermore, we also conduct human evaluation on this subset. We invite three graduate students specializing in NLP to participate in the evaluation. The human evaluation follows the same methodology used in the GPT evaluations. Please refer to Appendix F for more details.
@@ 4.2 Main Experiments
Tables 3(a) and 3(b) present the Rouge-L, GPT, and human evaluation scores respectively, for the instruction generalization. For the GPT-4 and human evaluation, each model is compared with our RoleGPT to determine a win rate, signifying the frequency at which a model is favored over RoleGPT. We observe that RoleLLaMA shows a considerable enhancement in role-playing performance compared to instruction-tuned baselines, the vanilla base model, and the previous role-playing models, in terms of speaking style imitation (CUS), response accuracy (RAW), and role-specific knowledge (SPE). Moreover, LLaMA-script directly tuned on original conversation data in scripts, even underperforms base LLaMA, which shows the imperative need for denoising of script data. The comparison between Alpaca/Vicuna and LLaMA indicates the efficacy of general-purpose instruction tuning in role-playing. However, the discrepancy between RoleLLaMA and Alpaca/Vicuna underscores the necessity to enhance their role-playing capabilities beyond mere general-purpose instruction-following. Besides, despite RoleLLaMA trailing behind RoleGPT in terms of speaking style imitation (CUS) and response accuracy (RAW)161616Note that drawing a direct comparison between RoleLLaMA and RoleGPT may not be equitable, given that RoleGPT is closed-source and the specifics of its model size and instruction tuning data remain undisclosed., it surpasses RoleGPT in role-specific knowledge, indicating the efficacy of our Context-Instruct technique.
@@
In Table 5(a) and Table 5(b), considering 10 held-out unseen roles, we observe that RoleLLaMA exhibits performance increase in imitating speaking style (CUS) and maintaining accuracy (RAW). However, for role-specific knowledge (SPE), there is not a noticeable improvement over the baselines, which is intuitive, since models with no prior knowledge of an unseen role would naturally exhibit confusion regarding it.
@@ 4.3 Ablation Study
| Method | SPE |
| RoleLLaMA (w/o c-inst) | 21.4 |
| RoleLLaMA (reaug, w/o c-inst) | 19.1 |
| RoleLLaMA (w/ c-inst) | 38.1 |
Effect of Context-Instruct.
To inject role-specific knowledge, a typical approach is to employ retrieval-augmentation role customization strategy (c.f., §2.2 & §4.3). In Table 6171717To ensure the relevance of the signals retrieved from the profiles, we only experiment on role-specific instructions., we explore the effect of Context-Instruct to inject role-specific knowledge, comparing it with retrieval augmentation and the absence of both techniques. We observe that Context-Instruct has the capability to substantially augment the models’ role-specific knowledge, and retrieval augmentation may lead to distraction and a lack of robustness due to the noisy nature of the retrieval source discussed in §4.3.
@@
| RoleGPT (zsp) | 9.3 | 2.4 |
| RoleGPT (fsp) | 29.8 | 1.9 |
| RoleGPT (fsd) | 63.3 | 1.5 |
@@
We employ GPT-4 to evaluate three role prompting methods in §2.2, namely, zero-shot prompt engineering (zsp), few-shot prompt engineering (fsp), and few-shot dialogue engineering (fsd). The evaluation is based on RoleBench-specific-zh in Table 7. The Win Rate is the frequency with which a method is ranked first by GPT-4 among the three and Avg. Ranking refers to the average ranking position of the method. We observe that few-shot dialogue engineering significantly outperforms few-shot prompt engineering, and both few-shot approaches excel over the zero-shot approach. These findings underscore the significance of dialogue-mode input formatting and retrieval augmentation for GPT models.
| Method | reaug | sys |
| RoleLLaMA | 36.7 | 38.1 |
| RoleGLM | 25.3 | 34.1 |
| RoleGPT | 23.6 | 23.4 |
@@
We compare two role customization methods: zero-shot system instruction (sys, §2.4) v.s. few-shot retrieval augmentation (reaug). Retrieval-augmentation-based role customization is akin to the way mentioned in §2.2, we insert retrieved dialogue pairs from profiles into the input, formatted based on the chat markup language of RoleLLaMA and RoleGLM. During both fine-tuning and inference, the input combines system instruction and the retrieved in-context demonstrations. Table 8 shows the superior performance of the system-instruction-based approach over the retrieval-augmentation-based approach for RoleLLaMA and RoleGLM, leading to a higher context efficiency. We suppose that the retrieved in-context examples from profiles are noisy and sparse, which leads to the distraction of the relatively small LLMs (e.g., RoleGLM and RoleLLaMA) and degrades their performance. While larger models such as RoleGPT exhibit greater robustness to noise and sparse information, leading to performance invariability and even performance increase (Li et al., 2022; Shi et al., 2023).
@@
In Figure 2, we analyze the scaling law of role-playing of RoleLLaMA with different sizes (i.e., 7B, 13B, 33B), and observe that a larger model leads to better results for role-playing.
We refer readers to Appendix G for more experimental analysis such as data mixing strategies, effect of Context-Instruct for role generalization, as well as cross-lingual and cross-cultural role-playing model performance.
5 Related Work
Role-Playing. Recent advances in the LLM community have showcased the potential of LLM customization and role-playing (Wei et al., 2023; Shanahan et al., 2023; Li et al., 2023a; Salemi et al., 2023; Maas et al., 2023; Li et al., 2023b; Chen et al., 2023a; Park et al., 2023). Playing specific roles enables LLMs to become more vivid (Li et al., 2023a), interactive (Maas et al., 2023; Wang et al., 2023b), personalized (Salemi et al., 2023; Halder, 2023), and capable of addressing complex tasks (Li et al., 2023b; Chen et al., 2023a; Qian et al., 2023). For example, Character.AI showcases the advanced role-playing capabilities inherent to LLMs. However, Character.AI is closed-source and doesn’t offer API access. Moreover, open-source LLMs still lag significantly behind SOTA closed-source LLMs such as ChatGPT and GPT-4 in terms of role-playing capability.
Data Augmentation for Instruction Tuning. Instruction tuning (aka., supervised fine-tuning, SFT) aims to enhance LLMs’ ability to follow instructions (Mishra et al., 2022; Longpre et al., 2023; Ouyang et al., 2022; Wang et al., 2023b; Zhang et al., 2023b). One of the main challenges for SFT is to obtain high-quality and diverse instruction data (Zhou et al., 2023; Wang et al., 2022a; Zhang et al., 2023a). However, manually annotating high-quality instruction data is costly. Fortunately, current SOTA LLMs such as GPT-4 (OpenAI, 2023) are increasingly considered superior to human annotators in various tasks and domains (Gilardi et al., 2023; Ding et al., 2023), in terms of both data quality and annotation cost. Thus, employing these LLMs for data augmentation is becoming an increasingly standard practice (Honovich et al., 2022; Taori et al., 2023; Xu et al., 2023; Ding et al., 2023; Qin et al., 2023b; Wang et al., 2023a).
@@ 6 Conclusion
We introduce RoleLLM to benchmark, elicit, and enhance role-playing in LLMs. We construct 100 role profiles, propose RoleGPT for speaking style imitation, and Context-Instruct for role-specific knowledge extraction. Through these methods, we create RoleBench, the first fine-grained benchmark and open-source instruction tuning data for role-playing. We obtain RoleLLaMA and RoleGLM by role-conditioned instruction tuning on RoleBench, which show strong role-playing abilities and are even comparable with RoleGPT (using GPT-4).
Limitations
Our framework is specifically designed for single-turn question-answering settings, which limits its applicability in multi-turn dialogue scenarios. Furthermore, our research and resource releases have been confined to English and Chinese languages. In future work, we plan to incorporate additional languages, with a focus on including more low-resource languages. Additionally, our dataset is currently limited to a selection of publicly available scripts. We see significant potential for expanding our dataset to encompass a wider array of scripts and characters.
Ethics Statement
Since role-playing may lead to jailbreaking (Liu et al., 2023), misuse of RoleLLM could generate sensitive or harmful content, reflecting the violent, explicit, or discriminatory traits of certain characters. This work and its assets are strictly for research purposes, opposing any malicious usage. It’s recommended to employ moderation and filtering mechanisms to curb adverse content dissemination. The responsibility lies with researchers and users to ensure ethical application of RoleLLM. We suggest the following measures to ensure a more responsible utilization of role-playing models:
1.
Implement Advanced Moderation Tools: Develop and integrate state-of-the-art content moderation tools that can automatically detect and filter out sensitive, harmful, or inappropriate content generated by RoleLLM.
2.
Bias Detection and Mitigation: Continuously work on detecting and mitigating biases in RoleLLM’s outputs, ensuring fairness and minimizing the perpetuation of stereotypes or discrimination.
3.
Transparency in Development and Use: Maintain a high level of transparency about the design, capabilities, and limitations of RoleLLM, including clear communication about the model’s decision-making processes.
4.
Feedback Mechanism: Implement a robust feedback system that allows users to report ethical concerns and problematic outputs, facilitating continuous improvement and accountability.
Appendix A Demonstrations
A.1 Non-Cherry-Picked Role-Playing Examples